Solve the ff problem.


3. What pressure is required to reduce the volume of a sample of air from 1.0 L to 0.0250

L? The original pressure on the sample is 98 kPa, the amount and temperature of the

sample remain fixed.


4. A 2450 ml sample of a gas at 285°C and 620 torr is compressed to a volume of 850 ml at

150°C. What is the new pressure of the gas?




1
Expert's answer
2021-04-08T12:08:33-0400

3. According to the combined gas law:

p1V1 = p2V2

where p1 - initial pressure, V1 - initial volume, p2 - final pressure, V2 - final volume.

From here:

p2 = p1V1 / V2 = 98 kPa × 1.0 L / 0.0250 L = 3920 kPa


Answer: 3920 kPa


4. According to the combined gas law:

p1V1/T1 = p2V2/T2

where p1 - initial pressure, V1 - initial volume, T1 - initial temperature (T1 = 285°C = 558.15 K) , p2 - final pressure, V2 - final volume, T2 - final temperature (T2 = 150°C = 423.15 K).

From here:

p2 = p1V1T2 / T1V2 = (620 torr × 2450 ml × 423.15 K) / (558.15 k × 850 ml) = 1354 torr


Answer: 1354 torr